The solubility of iron (II) hydroxide, Fe(OH)2, is 1.43 x10–3 gram per litre at 25 0C.

(a) Write a balanced equation for the solubility equilibrium.

(b) Write the expression for the solubility product constant, Ksp, and calculate its value.

(c) Calculate the pH of a saturated solution of Fe(OH)2 at 25 0C.

(d) A 50.0 millilitre sample of 3.00x10–3 molar FeSO4 solution is added to 50.0 millilitres of 4.00x10–6 molar NaOH solution. Does a precipitate of Fe(OH)2 form? Explain and show calculations to support your answer.
Answer #1

Dear Student, Kindly refer to the complete solution. Feel free to ask if you face difficulty in understanding any step of the solution attached so that we can assist you in a better way.

Dease Student, Given? - solubility of felony = 1643x108g/L at 25°c @ Balanced equation for solubility equilibciom, fe(OH) (s)③ [OHO ]= 3.18 x 105m. por= -log Cono = -log (3-18x105) = 4.50 pH = 14-por ( pH + por=14) PH = 14 - 4.50 = 9.50 @ when we are= (0.0015 M) x (2.00x1cm) = 6.00 x 1015 m3 Remember when kep> Qsp (Noppt) and when kep< Osp (ppt fornration) Kep = 1.61x1014